# AV-Link: Temporally-Aligned Diffusion Features for Cross-Modal Audio-Video Generation

### v1

@@ 4 Experiments
Diff-Foley [52] 11.00 28.71 0.06 7.88 0.115 0.121 0.14
S&H  [85] 29.22 66.51 0.025 2.09 0.179 0.189 0.128
FoleyCrafter [97] 4.62 18.66 0.080 9.10 0.204 0.215 0.166
Ours (VGGSounds) 2.02 13.68 0.129 10.06 0.203 0.214 0.405
Ours-Joint 2.19 13.15 0.122 9.24 0.205 0.215 0.409
Ours 1.58 14.17 0.137 9.93 0.207 0.223 0.531
S&H  [85] 10.72 25.44 0.186 6.06 0.30 0.318 0.08
FoleyCrafter [97] 2.99 18.05 0.212 10.88 0.208 0.219 0.179
@@
Ours 1.33 11.99 0.228 12.40 0.214 0.226 0.54
Table 1: Comparison of AV-Link against prior work on the V2A task on the VGGSounds [9] benchmark. We report a variant of our method with Fusion Blocks trained on VGGSounds only.
@@ 4.2 Evaluation Protocol
We separately evaluate our model on the V2A and A2V tasks. We consider a conditioning video or audio input from the test set and sample a corresponding audio or video using our V2A or A2V model respectively. As prior work does not consistently adopt conditioning text prompts at inference [52, 97, 62], we report results for both settings. In all experiments, we use 8 fusion blocks (186M parameters), a learning rate of 0.0001, and a batch size of 128 on 8 A100 GPUs for the ablation, increasing to 256 on 16 A100 GPUs for final experiments.
Baselines. For the task of V2A, we compare with Diff-foley [52], FoleyCrafter [97], Seeing and Hearing [85] (S&H), and Movie Gen [62]. Diff-foley conditions an audio generator on CAVP, an audio-aligned video feature extractor by concatenating it with text embeddings. FoleyCrafter adapts a pre-trained audio generator for A2V using a semantic adapter with frame-based CLIP features [63] and trains a temporal adapter that aligns audio based on an onset mask predicted from the video, inspired by ControlNet. S&H uses a training-free approach that optimizes cosine similarity in ImageBind [21] between video and generated audio. Movie Gen proposed to concatenate MetaClip [86] frame-wise features with audio tokens. For A2V, we compare with TempoToken [93], which optimizes latent tokens on a pre-trained video generator using BEATs [10] audio features.
@@ 4.3 Results
Video-to-Audio We present quantitative results comparing our method with baseline approaches on the V2A task in Tab. 1. Our method achieves the highest scores across all metrics, except for IB-AI and IB-AV, where S&H reports superior values. This is expected, as S&H explicitly optimizes for these metrics as part of its approach, making these scores less indicative of overall quality. Notably, our method demonstrates significantly better temporal alignment, with an Onset Accuracy of 0.53 compared to the top baseline score of 0.17. Such improvements are evident with and without providing audio text descriptions. For a more balanced comparison, we also evaluate our model trained only on VGGSounds, showing that it still substantially outperforms baselines, highlighting the benefit of using temporally-aligned diffusion features. This advantage is further confirmed by our user studies in Tab. 4, where AV-Link is consistently preferred, particularly for temporal alignment. Additionally, while Movie Gen achieves better sound quality due to its larger model (13B parameters), our method’s audio-video temporal alignment is preferred 63.6% of the time. We also illustrate our model’s strong temporal alignment in Fig. 4 and on the Website, where we perform inference on videos requiring precise temporal audio correspondence such as ‘bouncing’ and ‘tapping’ sounds. While baselines capture semantic content, they fall short in synchronizing the precise onset of these events.
Audio-to-Video On the A2V task, our method surpasses TempoTokens [77] in both video quality and audio-video alignment, as shown in Tab. 3. User studies in Tab. 4 also indicate that our approach is preferred for generation quality, semantic, and temporal alignment, with a preference greater than $74.8\%$ across all settings. Additionally, qualitative results in  Fig. 5 and on the Website demonstrate that our method produces videos that align closely with the audio signal, both semantically and temporally. Our generated video events (e.g., explosions) occur precisely in sync with corresponding audio events, showcasing the strength of our AV-Link in maintaining temporal coherence.
4.4 Ablation studies
Can V2A and A2V models share the same parameters? We jointly train V2A and A2V models by sharing the Fusion Block parameters between the two tasks. As reported in Tab. 1 and Tab. 3. Using a separate set of parameters contributes to a marginal but consistent improvement over parameter sharing.
Can both modalities be generated at the same time? Fig. 3 suggests a key issue with joint video-audio diffusion: Audio often depends on element that becomes only visible at the end of the sampling trajectory. For instance, accurately generating the sound of a gunshot requires high-frequency visual details, such as the smoke of an explosion, Similarly, sounds like thunder and explosions only become distant at the later sampler steps, making jointly denoising the two modalities challenging. This may explain the lower performance of current methods in this task [67, 85, 35], and suggest that audio-video generation benefits from factorization into video generation followed by V2A and vice-versa.
Which layers yield the best conditioning features? In Tab. 2, we ablate over four different designs where the multimodal Fusion Blocks are placed at early video DiT layers, in the middle, at the end, or regularly interleaved with the video DiT blocks. Interleaving the Fusion Blocks outperforms the other variants, indicating that feature diversity and regular modality alignment benefit the model.
Are Fusion Blocks necessary? We evaluate several methodologies for injecting conditioning features into the model for which we show results in the Feature injection section of Tab. 2. Concat. to text follows Diff-foley [52] in concatenating conditioning features to the text embeddings. Direct alignment follows Movie Gen [62] in repeating the visual features to match the number of audio frames and sum them together. Symm. cross attention replaces the symmetric design of Fusion Blocks based on self attention with two separate blocks based on cross attention: the first for injecting features into the audio backbone, the second for the video. The Fusion Block design consistently outperforms the baselines. Additionally, we explore using video backbone as a static feature extractor without reinjecting the output back to the subsequent video DiT blocks. While this approach allows for caching the video features and thus improves sampling efficiency, it significantly outperforms our method that progressively refines the conditioning features through reinjecting them back to subsequent DiT blocks.
How do diffusion features compare to pretrained representations? We experimented with various pretrained feature extractors in place of frozen generator activations as inputs to the Fusion Blocks. As shown in Tab. 2 diffusion features outperform all representations on the Movie Gen benchmark. On VGGSounds it shows the strongest temporal alignment as expressed by Onset ACC. User studies against baselines employing such representations (see Tab. 4) confirm improved temporal and semantic alignment.
TempoTokens [93] 103.09 2406.60 0.112 0.114
Ours-Joint 41.07 416.17 0.131 0.144
Ours 34.00 352.87 0.165 0.171
TempoTokens [93] 76.28 1247.70 0.167 0.173
Ours-Joint 32.89 297.75 0.193 0.206
Ours 32.90 228.68 0.206 0.210
Table 3: Comparison to baselines on the A2V task on VGGSounds [9].
@@ 4.4 Ablation studies
[…]
@@ 5 Conclusion
We present AV-Link, a unified approach for Video-to-Audio and Audio-to-Video generation. It leverages time-aligned activations from pretrained flow models, bypassing the use of specialized feature extractors. To enable this, we introduce a Fusion Block that allows for bidirectional information exchange between frozen generators for each modality through the use of temporally-aligned self attention and symmetric feature reinjection. Extensive ablations and comparisons show performance improvements over all metrics, with particular regard to temporal alignment.
@@ Appendix A Limitations
Our base video backbone is a low-resolution and low-fps RGB model. While the model achieves state-of-the-art V2A and A2V performance, leveraging a high-resolution latent video model may further improve performance. Exploring model scaling to further improve feature quality is an exciting avenue for future work. Additionally, the feature reinjection alters activations in the conditioning modality generator. While beneficial, it introduces additional compute at each sampling step as caching prior activations becomes infeasible. Step distillation techniques reduce this effect by reducing the number of sampling steps and constitute an orthogonal line of work.
@@ Appendix B Architecture Details
For the audio model, we employ a latent model with 24 DiT blocks with hidden dimension of 1024 for the patches. We use 16 attention heads for each self-attention and cross-attention operation, and an MLP hidden size of 4096. We follow [23] for encoding the audio and converting the generated Mel-spectrograms to waveform. Both models have 576M trainable parameters.
@@ Appendix C Training and Inference Details
This section presents additional details on training and inference, and discusses training and inference time.
@@ C.1 Training Details
For all training phases, we train our models using the AdamW optimizer with a learning rate of 3e-4, beta factors of 0.9 and 0.99, epsilon of 1e-8, a weight decay of 0.01 and a 10,000 steps warmup.
The base video model is trained for 250,000 steps on automatically-captioned internal dataset with a total batch size of 512 on 16 A100 GPUs. The base audio models is trained for 100,000 with a total batch size of 1024 on 8 A100 GPUs. We drop text condition 10% of the times to enable classifier free guidance (CFG) [28] during inference.
The fusion blocks are trained for 50,000 steps on 16 A100 GPUs with a batch size of 256. Ablation experiments are trained on 8 A100s with a total batch size 128. We drop the generated modality text prompt (e.g audio text prompt in V2A task) 50% of the time and the conditioning modality text prompt (e.g video text prompt in V2A task) 20% of the time. For all of our experiments, both the audio and video backbone are kept frozen unless otherwise specified.
@@ C.2 Inference details
We perform inference starting from pure gaussian noise for the modality to generate and use the model’s velocity estimates together with an Euler sampler to progressively transform the noise to the clean generated sample. We found using classifier free guidance on the conditioning modality to be instrumental towards obtaining good multimodal alignment. When conditioning on more than one modality (e.g. video and audio text prompt), we drop both conditions simultaneously to compute the unconditional signal. In all of our experiments, we use 64 sampling steps and a CFG weight of 5.0. While all of our generated video results are evaluated using the native 36x64 resolution, the generated videos displayed on the website have been upsampled from 36x64 to 144x256 spatial resolution using an internal upsampler for demo purposes.
@@ C.3 Training and Inference Time
Inference. We measure a throughput of 27.85s per sample using batch size of one on an A100 to perform 64 sampling steps for both the A2V and V2A tasks. This is a limitation of our method in the V2A task where the it results in a slower sampling time compared with previous approaches [52, 97]. We make, however, two important considerations. First, a major use case for V2A is sonification of generated soundless videos. Such generation, when performed by state-of-the-art large-scale text-to-video generators usually takes several minutes. Second, since our method relies on a flow model, distillation methods similar to the ones adopted by previous approaches [82] can significantly improve inference time by performing sampling in few steps. We regard this direction as as an interesting avenue for future work which is orthogonal to ours.
@@ Appendix D Additional Evaluation Results and Details
This section presents additional evaluation results an details. We highly encourage the reader to experience the generated audios and videos on the Website.
@@ D.2 Baselines Selection
Below, we include details on baselines selection and inference procedures.
@@
  • Frieren: Since the code was released 6 days before the deadline, we only included qualitative comparison with Frieren on the Website. We use the Frieren (reflow) model, 64 sampling steps, CFG of 5.0, and the other recommended hyper-parameters for inference.
@@
  • Seeing and Hearing: We use the official code for V2A and we follow the default sampling parameters. We exclude Seeing and Hearing qualitative comparison without prompt as it produces barely audiable sounds in this setting.
@@
Audio-to-Video. To the best of the authors’ knowledge, TempoToken [93] is the only in-the-wild A2V baseline with publicly available code. We exclude Seeing and Hearing [85] A2V results as their code is not available for this task. Additionally, we exclude joint audio-video generation methods such as MMDiffusion [67], as they were trained on the very limited Landscapes Videos dataset [42], which contains only 928 videos and lacks generalization beyond landscape scenarios. We also exclude sound-guided image animation methods [95, 42], as they address a fundamentally different task.
We use the offical implementation of TempoToken with default parameters to generate 2 seconds videos. For the quntitative comparison, we crop our generated videos to 2 seconds for a fair comparison with TempoToken.
D.3 Additional V2A results
Fig. 7 shows additional V2A results comparing our method to baselines. To better showcase the capabilities of AV-Link, we record a series of in-the-wild videos that require a high degree of temporal alignment for the audio modality and run inference for all baselines. Our method produces highly aligned audio results that capture the audio semantic entailed by the visual modality, while baselines produce degraded results. We attribute this phenomenon to the lack of access to visual features that are precisely aligned to the video content. On the contrary, the use of activations from the video generation backbone allows AV-Link to produce precise alignment in this scenario.
D.4 Additional A2V results
In Fig. 8 we show additional visualizations of A2V results produced by AV-Link in the case where textual descriptions for the visual modality are given, and refer the reader to the Website for A2V results produced without textual description. AV-Link can generate visual content that is well-aligned to the audio modality.

### v2

@@ 4 Experiments
[…]
$\text{FoleyCrafter}^{*}$ [104] 4.62 18.66 - 9.10 0.204 0.215 0.285
$\text{Frieren}^{\dagger}$ [89] 3.52 16.50 - 7.78 0.178 0.184 0.301
[…]
Ours (VGGSounds) 2.02 13.68 - 10.06 0.203 0.214 0.405
Ours-Joint 2.19 13.15 - 9.24 0.205 0.215 0.409
Ours 1.58 14.17 - 9.93 0.207 0.223 0.531
[…]
$\text{FoleyCrafter}^{*}$ [104] 2.99 18.05 0.212 10.88 0.208 0.219 0.307
@@
Ours 1.33 11.99 0.228 12.40 0.214 0.226 0.540
[…]
@@ 4.2 Evaluation Protocol
We separately evaluate our model on the V2A and A2V tasks. We consider a conditioning video or audio input from the test set and sample a corresponding audio or video using our V2A or A2V model respectively. As prior work does not consistently adopt conditioning text prompts at inference [57, 104, 68], we report results for both settings. In all experiments, we use 8 fusion blocks (186M parameters), a learning rate of $41e\!-\!$ 4, and a batch size of 128 on 8 A100 GPUs for the ablation, increasing to 256 on 16 A100 GPUs for final experiments.
Baselines. For the task of V2A, we compare with Diff-foley [57], FoleyCrafter [104], Seeing and Hearing [92] (S&H), Frieren [89], V2A Mapper [86] and Movie Gen [68]. Diff-foley and V2A-Mapper conditions an audio generator on CAVP and CLIP, respectively by concatenating it with text embeddings. FoleyCrafter adapts a pre-trained audio generator for A2V using a semantic adapter with frame-based CLIP features [69] and trains a temporal adapter that aligns audio based on an onset mask predicted from the video, inspired by ControlNet. S&H uses a training-free approach that optimizes cosine similarity in ImageBind [25] between video and generated audio. Movie Gen and Frieren proposed to concatenate frame-wise features with audio tokens. For A2V, we compare with TempoToken [100], which optimizes latent tokens on a pre-trained video generator using BEATs [11] audio features.
TempoTokens [100] 81.43 1489.35 0.112 0.114
Ours-Joint 51.63 297.24 0.112 0.126
Ours 39.10 266.93 0.160 0.168
TempoTokens [100] 77.87 1275.50 0.167 0.173
Ours-Joint 40.72 206.30 0.185 0.192
Ours 37.62 158.71 0.191 0.201
Table 3: Comparison to baselines on the A2V task on VGGSounds [10].
@@ 4.3 Results
Video-to-Audio We compare AV-Link with baseline approaches on the V2A task in Tab. 1. Our method achieves the highest scores across all metrics, except for IB-AI and IB-AV, where S&H reports superior values. This is expected, as S&H explicitly optimizes for these metrics as part of its approach, making these scores less indicative of overall quality. Notably, our method demonstrates significantly better temporal alignment, with an Onset Accuracy of 0.53 compared to the top baseline score of 0.3. Such improvements are evident with and without providing audio text descriptions. For a more balanced comparison, we also evaluate our model trained only on VGGSounds, showing that it still substantially outperforms baselines. This advantage is further confirmed by our user studies in Tab. 4, where AV-Link is consistently preferred, particularly for temporal alignment. Additionally, while Movie Gen achieves better sound quality due to its larger model (13B parameters), our method’s audio-video temporal alignment is preferred 63.6% of the time. We also illustrate our model’s strong temporal alignment in Fig. 4 and on the Website, where we perform inference on videos requiring precise temporal audio correspondence such as ‘bouncing’ and ‘tapping’ sounds. While baselines capture semantic content, they fall short in synchronizing the precise onset of these events.
@@
Audio-to-Video On the A2V task, our method surpasses TempoTokens [83] in both video quality and audio-video alignment, as shown in Tab. 3. User studies in Tab. 4 also indicate that our approach is preferred for generation quality, semantic, and temporal alignment, with a preference greater than $73.9\%$ across all settings. Additionally, qualitative results in  Fig. 5 and on the Website demonstrate that our method produces videos that align closely with the audio signal, both semantically and temporally. Our generated video events (e.g., explosions) occur precisely in sync with corresponding audio events, showcasing the strength of our AV-Link in maintaining temporal coherence.
4.4 Ablation studies
Can V2A and A2V models share the same parameters? We jointly train V2A and A2V models by sharing Fusion Block parameters of the two tasks. As reported in Tab. 1 and Tab. 3, using separate parameters contributes to a marginal but consistent improvement over parameter sharing.
Can both modalities be generated at the same time? Fig. 3 suggests a key issue with joint video-audio diffusion: Audio often depends on element that becomes only visible at the end of the sampling trajectory. For instance, accurately generating the sound of a gunshot requires high-frequency visual details, such as the smoke of an explosion, Similarly, sounds like thunder and explosions only become distant at the later sampler steps, making jointly denoising the two modalities challenging. This may explain the lower performance of current methods in this task [73, 92, 87], and suggest that audio-video generation benefits from factorization into video generation followed by V2A and vice-versa. Indeed, when we train AV-Link for joint audio-video generation, we observed a substantial drop in performance.
Figure 5: Qualitative results of A2V generation. Our model generates semantically and temporally aligned content capturing temporal events implied by the audio modality such as explosions and printing. See the Website for additional results.
Are Fusion Blocks necessary? We evaluate several methodologies for injecting conditioning features into the model for which we show results in the Feature injection section of Tab. 2. Concat. to text follows Diff-foley [57] in concatenating conditioning features to the text embeddings. Direct alignment follows Movie Gen [68] in repeating the visual features to match the number of audio frames and sum them together. Symm. cross attention replaces the symmetric design of Fusion Blocks based on self attention with two separate blocks based on cross attention: the first for injecting features into the audio backbone, the second for the video. The Fusion Block design consistently outperforms the baselines. Additionally, we explore using video backbone as a static feature extractor without reinjecting the output back to the subsequent video DiT blocks. While this approach allows for caching the video features and thus improves sampling efficiency, it significantly underperforms our method that progressively refines the conditioning features by reinjecting them back to subsequent DiT blocks.
How do diffusion features compare to pretrained representations? We experimented with various pretrained feature extractors in place of frozen generator activations as inputs to the Fusion Blocks. As shown in Tab. 2, diffusion features outperform all representations on the Movie Gen benchmark. On VGGSounds it shows the strongest temporal alignment as expressed by Onset ACC. User studies against baselines employing such representations (see Tab. 4) confirm improved temporal and semantic alignment.
Which layers yield the best conditioning features? In Tab. 2, we ablate over four different designs where the multimodal Fusion Blocks are placed at early video DiT layers, in the middle, in the end, or regularly interleaved with the video DiT blocks. Interleaving the Fusion Blocks outperforms the other variants, indicating that feature diversity and regular modality alignment benefit the model.
@@ 5 Conclusion
We present AV-Link, a unified approach for V2A and A2V generation. It leverages time-aligned activations from pretrained flow models, bypassing the use of specialized feature extractors. To enable this, we introduce a Fusion Block that allows for bidirectional information exchange between frozen generators for each modality through the use of temporally-aligned self attention and symmetric feature reinjection. Extensive ablations and comparisons show performance improvements over all metrics, with particular regard to temporal alignment.
@@ Appendix A Limitations
Our base video backbone is a low-resolution and low-fps RGB model coupled with a high-resolution upsampler model. While the model achieves state-of-the-art V2A and A2V performance, leveraging a large high-resolution latent video model may further improve performance. Exploring model scaling to further improve feature quality is an exciting avenue for future work. Additionally, the feature reinjection alters activations in the conditioning modality generator. While beneficial, it introduces additional compute at each sampling step as caching prior activations becomes infeasible. Step distillation techniques reduce this effect by reducing the number of sampling steps and constitute an orthogonal line of work.
@@ Appendix B Architecture Details
For the audio model, we employ a latent model with 24 DiT blocks with hidden dimension of 1024 for the patches. We use 16 attention heads for each self-attention and cross-attention operation and an MLP hidden size of 4096. We follow [27] for encoding the audio and converting the generated Mel-spectrograms to waveform. Both models have 576M trainable parameters.
@@ Appendix C Training and Inference Details
This section presents additional details on training and inference and discusses training and inference time.
@@ C.1 Training Details
For all training phases, we train our models using the AdamW optimizer with a learning rate of 3e-4, beta factors of 0.9 and 0.99, epsilon of 1e-8, a weight decay of 0.01, and a 10,000-step warmup.
The base video model is trained for 250,000 steps on an automatically-captioned internal dataset with a total batch size of 512 on 16 A100 GPUs. The base audio models are trained for 100,000 with a total batch size of 1024 on 8 A100 GPUs. We drop text condition 10% of the time to enable classifier-free guidance (CFG) [31] during inference.
The fusion blocks are trained for 50,000 steps on 16 A100 GPUs with a batch size of 256. Ablation experiments are trained on 8 A100s with a total batch size of 128. We drop the generated modality text prompt (e.g. audio text prompt in V2A task) 50% of the time and the conditioning modality text prompt (e.g. video text prompt in V2A task) 20% of the time. For all of our experiments, both the audio and video backbone are kept frozen unless otherwise specified.
@@ C.2 Inference details
We perform inference starting from pure Gaussian noise for the modality to generate and use the model’s velocity estimates together with an Euler sampler to progressively transform the noise to the clean generated sample. We found using classifier-free guidance on the conditioning modality to be instrumental in obtaining good multimodal alignment. When conditioning on more than one modality (e.g. video and audio text prompt), we drop both conditions simultaneously to compute the unconditional signal. In all of our experiments, we used 64 sampling steps and a CFG weight of 5.0. All of our generated video results are evaluated at the 512 x 288 pixels resolution.
@@ C.3 Training and Inference Time
Inference. We measure a throughput of 27.85s per sample using a batch size of one on an A100 to perform 64 sampling steps for both the A2V and V2A tasks. This is a limitation of our method in the V2A task since it results in a slower sampling time compared with previous approaches [57, 104]. We make, however, two important considerations. First, a major use case for V2A is the sonification of generated soundless videos. Such generation, when performed by state-of-the-art large-scale text-to-video generators usually takes several minutes. Second, since our method relies on a flow model, distillation methods similar to the ones adopted by previous approaches [89] can significantly improve inference time by performing sampling in a few steps. We regard this direction as an interesting avenue for future work that is orthogonal to ours.
@@ Appendix D Additional Evaluation Results and Details
This section presents additional evaluation results and details. We highly encourage the reader to experience the generated audios and videos on the Website.
@@ D.2 Baselines Selection
Below, we include details on baseline selection and inference procedures.
@@
  • Frieren: We use the Frieren (reflow) model, 64 sampling steps, CFG of 5.0, and the other recommended hyper-parameters for inference.
@@
  • Seeing and Hearing: We use the official code for V2A and we follow the default sampling parameters. We exclude Seeing and Hearing qualitative comparison without prompt as it produces barely audible sounds in this setting.
@@
Audio-to-Video. To the best of the authors’ knowledge, TempoToken [100] is the only in-the-wild A2V baseline with publicly available code. We exclude Seeing and Hearing [92] A2V results as their code is not available for this task. Additionally, we exclude joint audio-video generation methods (see Appx. D.3) such as MMDiffusion [73], as they were trained on the very limited Landscapes Videos dataset [46], which contains only 928 videos and lacks generalization beyond landscape scenarios. We also exclude sound-guided image animation methods [103, 46], as they address a fundamentally different task.
We use the official implementation of TempoToken with default parameters to generate 2-second videos. For the quantitative comparison, we crop our generated videos to 2 seconds for a fair comparison with TempoToken.
D.3 Joint Audio-Video Generation Baselines
In this work, we aim to address the tasks of A2V and V2A generation within a single framework. Some methods for joint video-audio generation [28, 40, 97, 50, 81, 78] are capable of operating under this conditional setting.
However, due to the difficulties in jointly modeling the audio and video modalities, these methods are often trained on domain specific datasets: Landscapes [46] is composed of 928 videos of natural landscapes; AIST++ [48] contains 1020 clips (5.2 hours) of dancing human sequences; GreatestHits [64] is composed of 977 videos featuring a drumstick hitting objects in a scene; EPIC-SOUNDS [34] encompasses 117.6k clips (100 hours) of cooking-related actions; Monologues [40] features 19.1M clips of talking people. The use of narrow distribution datasets coupled with limited availability of source code and pretrained checkpoints limits the possibility of performing meaningful comparisons on the broad data distribution modeled by AV-Link.
In the following, we discuss the considered joint audio-video baseline methods. MMDiffusion [73] provides checkpoints for models trained on Landscapes [46] and AIST++ [48] datasets only. We find that the Landscapes checkpoint overfits to its training dataset distribution, frequently replicating training samples. By analyzing 500 generated videos using the provided checkpoint, we found that the generated videos achieved an average of 0.825 CLIP similarity between the generated videos and the top-matching video from the training dataset. We include in Fig. 8 examples of such overfitting. This prevents the method from operating on videos outside these domains and does not allow for an informative comparison. Ishii et al. [36] provides a checkpoint trained on the GreatestHits [64] dataset only. While trained also on the Landscapes [46] and VGGSound [10] datasets, no A2V or V2A results were reported. Kim et al. [40] report results on the Landscapes [46], AIST++ [48] and the Monologues (talking heads) datasets with no code publicly available. CMMD [97] report results on the AIST++ [48] and EPIC-SOUNDS [34] datasets and do not provide code.
D.4 Additional V2A results
Fig. 7 shows additional V2A results comparing our method to baselines. To better showcase the capabilities of AV-Link, we record a series of in-the-wild videos that require a high degree of temporal alignment for the audio modality and run inference for all baselines. Our method produces highly aligned audio results that capture the audio semantic entailed by the visual modality, while baselines produce degraded results. We attribute this phenomenon to the lack of access to visual features that are precisely aligned with the video content. On the contrary, the use of activations from the video generation backbone allows AV-Link to produce precise alignment in this scenario.